OpenStudy (kittiwitti1):

Need conversion units (don't do any problems for me) to check! a. "How many moles of oxygen are present in 1.39 • 10^22 molecules of water?" b. "How many atoms of hydrogen are in 88.5 grams of water?" c. "How many Ti atoms are contained in 12.80 grams of Ti?" d. "What is the mass of 3.25 mol of Pt?"

OpenStudy (kittiwitti1):

assumed conversions: a) \(\LARGE{\frac{1molO}{6.022\times10^{23}moleculesO}}\) b) \(\LARGE{\frac{1molH}{1.008gH}}\),\(\LARGE{\frac{1molH_{2}O}{2molH}}\),\(\LARGE{\frac{1molH}{6.022\times10^{23}moleculesH}}\)

OpenStudy (kittiwitti1):

c) \(\LARGE{\frac{1molTi}{47.867gTi}}\) d) \(\LARGE{\frac{195.084gPt}{1molPt}}\)

OpenStudy (kittiwitti1):

the last fraction in B should be reversed and have \(H_{2}O\) oops also it's \(\cancel{molecules}→atoms\) my bad

OpenStudy (sweetburger):

A. looks good. B. Your method might be correct, but it seems strange to do it that way. I would do it like this\[\frac{ 88.5gH_2O }{ 18.0g/molH_2O }\times \frac{ 2moles H }{ 1moleH_2O }\times \frac{ 6.022\times 10^\left( 23 \right)atomsH }{ 1 mole H }\]

OpenStudy (sweetburger):

C. is fine except that you must also multiply by \(6.022\times10^\left( 23 \right)\)

OpenStudy (kittiwitti1):

I corrected B though ... is it still wrong lol

OpenStudy (sweetburger):

Im comparing the answer I get from my process to yours. Just to see if they diverge. one sec

OpenStudy (kittiwitti1):

okay

OpenStudy (sweetburger):

Ok I think you may have a mistake in your process. The mistake is that when you divide 88.5 H2O by 1.008g H the rest of your calculation assumes that you end up having roughly 87moles of H which is not true. 88.5g of H2O has closer to 9.833 moles of H.

OpenStudy (kittiwitti1):

Alright

OpenStudy (sweetburger):

Even when you consider the division by 2. 87/2 = 43.5 moles of H still way too high.

OpenStudy (sweetburger):

And for d you did it perfect.

OpenStudy (kittiwitti1):

Wait... so for B I used the wrong designation (H instead of H2O) in the first fraction(s)?

OpenStudy (sweetburger):

Yes, you need to divide by 18.0grams of H2O not 1.008 grams of H.

OpenStudy (kittiwitti1):

Ahh, I see o: Alright thank you! ☺

OpenStudy (sweetburger):

Yea, np glad to help :)